OpenStudy (anonymous):

Write an equation of the line with y-intercept -6 that is parallel to the graph of x - 3y = 8.

OpenStudy (anonymous):

a.y=1/3x-6 b.3y=x-8 c.3y = x - 6 d.y = 4x - 6

OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –5 when x = 15, find y when x = –1

OpenStudy (texaschic101):

x - 3y = 8 -3y = -x + 8 y = 1/3x - 8/3 slope of this line is 1/3.....a parallel line will have same slope. so we want a line with slope of 1/3 and y intercept of -6. In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So your parallel line is : y = 1/3x - 6

OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = –48 when x = –8, find y when x = –3. a. y = 6x; –18 c. y = 6x; –21 b. y = 5x; –18 d. y = –6x; 18

OpenStudy (texaschic101):

y/x = k -5/15 = k -1/3 = k direct variation is y = -1/3x y = -1/3x --- when x = -1 y = -1/3(-1) y = 1/3 ========================= y/x = k -48/-8 = k 6 = k y = 6x --- when x = -3 y = 6(-3) y = -18

OpenStudy (anonymous):

a. y = 6x; –18 c. y = 6x; –21 b. y = 5x; –18 d. y = –6x; 18 which one is it? @texaschic101

OpenStudy (texaschic101):

answer A ... y = 6x : -18